### v2

# The thermal view of $f(R)$ cosmology

Abstract
A new thermal view of scalar-tensor gravity, in which general relativity is the zero-temperature state of gravity, is applied to the subclass of $f(R)$ gravity theories and, specifically, to spatially homogeneous and isotropic universes. Within the limits of application of the new thermal formalism, results on the convergence to Einstein cosmology (or lack thereof) are first obtained for general $f(R)$ theories, and then illustrated with power-law and Starobinsky $f(R)$ gravity.
I Introduction
General Relativity (GR) has been very successful in the regimes where it is tested with good precision, but it cannot be the ultimate theory of gravity. If gravity and quantum mechanics have to merge at the Planck scale, as it is widely believed, GR must be changed already at low energy. The smallest quantum corrections modify the Einstein-Hilbert action by raising the order of the field equations and introducing extra propagating degrees of freedom [1, 2]. Likewise, the low-energy limit of the bosonic string (the simplest string theory, which still incorporates essential features of string theories) does not reproduce GR but gives, instead, an $\omega=-1$ Brans-Dicke theory [3, 4] (where $\omega$ is the Brans-Dicke coupling constant introduced below).
The simplest way of adding a new degree of freedom to GR is by making Newton’s constant $G$ a field varying across space and time, as proposed in the original Brans-Dicke theory [5] containing a gravitational scalar field $\phi$ and an effective gravitational coupling $G_{\mathrm{eff}}\simeq 1/\phi$. Brans-Dicke theory was later generalized to “first generation” scalar-tensor gravity [6, 7, 8, 9], still containing only a single extra scalar field $\phi$ in addition to GR, but making this scalar massive by endowing it with a potential $V(\phi)$, plus allowing the coefficient of the kinetic term of $\phi$ in the action to become a function of $\phi$ [6, 7, 8, 9].
Enter cosmology: early universe inflation, although not proved, is widely regarded as a paradigm of modern cosmology. Starobinsky inflation [10], which is based on quadratic corrections to the GR Lagrangian, stands as the most viable scenario of inflation and does not require the introduction of an ad hoc inflaton field, since the quadratic corrections that modify GR produce the same dynamics.
[…]
In the last decade, the quest for the most general scalar-tensor gravity with equations of motion of only second order has led to rediscovering Horndeski gravity [17], and to further generalize it with Degenerate Higher Order Scalar-Tensor (DHOST) theories. There is now a vast literature on scalar-tensor gravity and cosmology and it is difficult to gain a comprehensive view of all their aspects and observational constraints at different scales and regimes across the universe. A recent proposal ([18, 19, 20, 21, 22, 23, 24, 25, 26], see also [27]) views scalar-tensor gravity as a thermally excited state of GR, which corresponds to the “zero-temperature” state of gravity. The approach of scalar-tensor gravity to GR occurring, for example, in the early universe [28, 29] (with its complications [30]), is seen as the analogue of heat conduction in an effective dissipative fluid associated with the scalar degree of freedom $\phi$. A new formalism describing these situations (under certain restrictions) is currently under development and is being tested on various scalar-tensor theories and their exact solutions. Here we continue this work, applying this thermal picture (described below) to $f(R)$ cosmology. In this context, the new thermal view of scalar-tensor gravity unifies several theoretical results that appeared in a rather fragmented literature over the years and predicts when, and under what conditions, $f(R)$ universes converge to, or depart from, GR ones. Two basic ideas of the thermal view of scalar-tensor gravity are: i) gravity is “hot” (i.e., deviates from GR) in strong gravity, in particular near spacetime singularities; ii) the expansion of 3-space “cools” gravity, making it approach GR analogously to heat conduction in a dissipative fluid. Although these basic ideas prove true in simplified situations, more involved regimes involve the presence of “heat sources” or “sinks” in the relevant equations, which complicate the basic picture. These more involved situations occur in $f(R)$ cosmology, as discussed in the following sections.
The new thermal formalism basically consists of an analogy between the dynamics of gravity and heat dissipation in a relativistic imperfect fluid. This analogy has limitations in its regime of applicability and, like all analogies, when using it one should not confuse its two sides, which describe physically different systems. However, since there are two sides to the analogy, one can use one of them (the most intuitive and better known one) to understand the other (less known) side, which is what we do here.
To begin, let us introduce scalar-tensor gravity and its new thermal description. We follow the notation of Ref. [31], using units in which the speed of light $c$ and Newton’s constant $G$ are unity, and the metric signature is +++{-}{+}{+}{+}. First-generation scalar-tensor gravity is described by the Jordan frame action [5, 6, 7, 8, 9]
$S_{\mathrm{ST}}=\int\frac{d^{4}x}{16\pi}\sqrt{-g}\left[\phi R-\frac{\omega}{\phi}\,\nabla^{c}\phi\nabla_{c}\phi-V(\phi)\right]+S^{\mathrm{(m)}}\,$, (1.1)
where $g$ is the determinant of the spacetime metric $g_{ab}$ with Ricci scalar $R$, $\phi>0$ is the Brans-Dicke-like scalar field (approximately corresponding to the inverse of the effective gravitational coupling $G_{\mathrm{eff}}\simeq 1/\phi$), $V(\phi)$ is the scalar field potential, the “Brans-Dicke coupling” $\omega(\phi)>-3/2$ to avoid $\phi$ being a phantom field, and $S^{\mathrm{(m)}}$ is the matter action. The Jordan frame field equations read
$R_{ab}-\frac{1}{2}\,g_{ab}R=\frac{8\pi}{\phi}\,T_{ab}^{\mathrm{(m)}}+T_{ab}^{(\phi)}\,$, (1.2)
$\,\,\,T_{ab}^{(\phi)}=\frac{\omega}{\phi^{2}}\left(\nabla_{a}\phi\nabla_{b}\phi-\frac{1}{2}\,g_{ab}\nabla^{c}\phi\nabla_{c}\phi\right)$
$\qquad\quad\quad+\frac{1}{\phi}\left(\nabla_{a}\nabla_{b}\phi-g_{ab}\Box\phi\right)-\frac{V}{2\phi}\,g_{ab}\,$,\, (1.3)
$(2\omega+3)\Box\phi=8\pi T^{\mathrm{(m)}}+\phi\,V^{\prime}-2V-\frac{d\omega}{d\phi}\,\nabla^{c}\phi\nabla_{c}\phi\,$,
where $R_{ab}$ is the Ricci tensor,
$T_{ab}^{\mathrm{(m)}}=\frac{-2}{\sqrt{-g}}\,\frac{\delta S^{\mathrm{(m)}}}{\delta g^{ab}}$ (1.5)
is the matter stress-energy tensor, $T^{(m)}\equiv g^{ab}T_{ab}^{\mathrm{(m)}}$, $\nabla_{a}$ is the covariant derivative operator of $g_{ab}$, $\Box\equiv g^{ab}\nabla_{a}\nabla_{b}$, and $T_{ab}^{(\phi)}$ is an effective stress-energy tensor of $\phi$.
Let us come now to the new thermal view of scalar-tensor gravity [18, 19, 20, 21, 22, 23, 24, 25, 26] (here we limit ourselves to “first-generation” scalar-tensor theories, which is all we need for $f(R)$ cosmology). The basic idea is to regard $T_{ab}^{(\phi)}$ in Eq. (1.3) as the effective stress-energy tensor of a fluid, which is possible when the gradient $\nabla^{a}\phi$ is timelike and future-oriented (a basic restriction of the formalism). Then, we define the effective fluid’s four-velocity
$u^{a}\equiv\frac{\nabla^{a}\phi}{\sqrt{-\nabla_{c}\phi\nabla^{c}\phi}}$ (1.6)
and $T_{ab}^{(\phi)}$ assumes the structure of a dissipative fluid stress-energy tensor [32, 33, 34]
$T_{ab}^{(\phi)}=\rho^{(\phi)}u_{a}u_{b}+P^{(\phi)}h_{ab}+\pi_{ab}^{(\phi)}+q_{a}^{(\phi)}u_{b}+q_{b}^{(\phi)}u_{a}\,$, (1.7)
where $h_{ab}\equiv g_{ab}+u_{a}u_{b}$ is the Riemannian metric on the 3-space experienced by observers comoving with the fluid,
$\rho^{(\phi)}=-\frac{\omega\,\nabla^{e}\phi\nabla_{e}\phi}{2\phi^{2}}+\frac{V}{2\phi}+\frac{1}{\phi}\left(\square\phi-\frac{\nabla^{a}\phi\nabla^{b}\phi\nabla_{a}\nabla_{b}\phi}{\nabla^{e}\phi\nabla_{e}\phi}\right)$
is an effective energy density,
$P^{(\phi)}=-\frac{\omega\,\nabla^{e}\phi\nabla_{e}\phi}{2\phi^{2}}-\frac{V}{2\phi}-\frac{1}{3\phi}\left(2\square\phi+\frac{\nabla^{a}\phi\nabla^{b}\phi\nabla_{b}\nabla_{a}\phi}{\nabla^{e}\phi\nabla_{e}\phi}\right)$ (1.9)
is an effective isotropic pressure,
$\pi_{ab}^{(\phi)}$ $=$ $\frac{1}{\phi\nabla^{e}\phi\nabla_{e}\phi}\left[\frac{1}{3}\left(\nabla_{a}\phi\nabla_{b}\phi-g_{ab}\nabla^{c}\phi\nabla_{c}\phi\right)\left(\square\phi-\frac{\nabla^{c}\phi\nabla^{d}\phi\nabla_{d}\nabla_{c}\phi}{\nabla^{e}\phi\nabla_{e}\phi}\right)\right$. (1.10)
$\left.+\nabla^{d}\phi\left(\nabla_{d}\phi\nabla_{a}\nabla_{b}\phi-\nabla_{b}\phi\nabla_{a}\nabla_{d}\phi-\nabla_{a}\phi\nabla_{d}\nabla_{b}\phi+\frac{\nabla_{a}\phi\nabla_{b}\phi\nabla^{c}\phi\nabla_{c}\nabla_{d}\phi}{\nabla^{e}\phi\nabla_{e}\phi}\right)\right]$
is the (purely spatial and trace-free) stress tensor, while
$q_{a}^{(\phi)}=\frac{\nabla^{c}\phi\nabla^{d}\phi}{\phi\left(-\nabla^{e}\phi\nabla_{e}\phi\right)^{3/2}}\,\Big(\nabla_{d}\phi\nabla_{c}\nabla_{a}\phi-\nabla_{a}\phi\nabla_{c}\nabla_{d}\phi\Big)$ (1.11)
is an effective heat flux density [18]. The second derivatives of $\phi$ endow this effective fluid with a dissipative character. This structure is common to all symmetric two-index tensors and there is no physics in it [35], but then a little miracle happens. Eckart’s theory of dissipative fluids [32, 33, 34] assumes three constitutive relations, the most important of which is the generalized Fourier law
$q_{a}=-{\cal K}h_{ab}\left(\nabla^{b}{\cal T}+{\cal T}\dot{u}^{b}\right)\,$, (1.12)
where $\cal K$} is the thermal conductivity of the fluid, $\cal T$} is its temperature, and $\dot{u}^{a}\equiv u^{c}\nabla_{c}u^{a}$ is the fluid four-acceleration. A direct computation [18] shows that $q_{a}^{(\phi)}$ is proportional to $\dot{u}_{a}$, which allows one to identify its coefficient with
${\cal K}{\cal T}=\frac{\sqrt{-\nabla^{c}\phi\nabla_{c}\phi}}{8\pi\phi}\,$, (1.13)
thus defining the product of the “effective thermal conductivity” and the “effective temperature of gravity”. GR is obtained for $\phi=$ const. and corresponds to zero temperature, the state of thermal equilibrium. Regimes in which the scalar degree of freedom is excited and propagates correspond to “warmer” states at ${\cal K}{\cal T}>0$ [19, 20, 21, 22, 23, 24, 25, 26]. A number of results have been obtained in the formalism, including the evolution equation for ${\cal K}{\cal T}$ [19, 20, 21, 22]
$\frac{d\left({\cal K}{\cal T}\right)}{d\tau}$ $=$ $8\pi\left({\cal K}{\cal T}\right)^{2}-\Theta\,{\cal K}{\cal T}+\frac{T^{\mathrm{(m)}}}{\left(2\omega+3\right)\phi}$
$+\frac{1}{8\pi\left(2\omega+3\right)}\left(V^{\prime}-\frac{2V}{\phi}-\frac{1}{\phi}\,\frac{d\omega}{d\phi}\,\nabla^{c}\phi\nabla_{c}\phi\right)\,$,
where $\tau$ is the proper time along the effective $\phi$-fluid lines and $\Theta\equiv\nabla_{c}u^{c}$ is its expansion scalar.
Gravity is “heated” (i.e., $d\left({\cal K}{\cal T}\right)/d\tau>0$) by positive terms in the right-hand side and is “cooled” (i.e., $d\left({\cal K}{\cal T}\right)/d\tau<0$) by negative ones.
The recent Ref. [25] advances the thermal picture with the study of scalar-tensor situations in which $\Box\phi=0$, in which case the $\left(\Theta,{\cal K}{\cal T}\right)$ plane (not a phase space) provides a convenient representation of the behavior of gravity. If the expansion scalar $\Theta$ is negative, or if the system begins at a point above the critical half-line $8\pi{\cal K}{\cal T}=\Theta>0$, gravity either “heats up” diverging away from GR, or else this line is crossed with horizontal tangent, after which gravity cools forever [25, 36].
If instead $\Theta>0$ and the system begins at a point below this critical half-line (i.e., with $8\pi{\cal K}{\cal T}<\Theta$), gravity can only “cool” and converge to GR [25]. This picture survives if conformal matter with $T^{(m)}=0$, but no potential, is added to the picture. Here we study vacuum $f(R)$ gravity, which unavoidably adds a potential for $\phi=f^{\prime}(R)$.
Section II reviews $f(R)$ gravity and cosmology, while the following sections apply the thermal view to it. Several exact solutions of $f(R)$ gravity are known (e.g., [37, 38, 39, 40, 41]) and we use some of them as examples.
II $f(R)$ gravity
We focus on a subclass of scalar-tensor gravities, i.e., metric $f(R)$ gravity described by the action
$S=\frac{1}{16\pi}\int d^{4}x\sqrt{-g}\,f(R)+S^{\mathrm{(m)}}\,$, (2.1)
where $f(R)$ is a non-linear function of the Ricci scalar. The fourth order field equations produced by varying the action (2.1) with respect to the inverse metric $g^{ab}$ are [14, 15, 16]
$f^{\prime}(R)R_{ab}-\frac{f(R)}{2}\,g_{ab}=8\pi\,T_{ab}^{\mathrm{(m)}}+\left(\nabla_{a}\nabla_{b}-g_{ab}\Box\right)f^{\prime}(R)\,$, (2.2)
where a prime denotes differentiation with respect to the argument. Tracing Eq. (2.2) yields
$\Box f^{\prime}+\frac{1}{3}\left[f^{\prime}(R)R-2f(R)\right]=\frac{8\pi}{3}\,T^{\mathrm{(m)}}\,$, (2.3)
which implies that $f^{\prime}(R)$ is a dynamical, propagating degree of freedom.
It is well known [14, 15, 16] that metric $f(R)$ gravity is equivalent to a Brans-Dicke theory with scalar field $\phi=f^{\prime}(R)$, Brans-Dicke parameter $\omega=0$, and scalar field potential
$V(\phi)=\phi R-f(R)\Bigg|_{f^{\prime}(R)=\phi}\,.$ (2.4)
In general, this potential cannot be written explicitly in terms of $\phi$.
In $f(R)$ gravity, it must be $f^{\prime}(R)>0$ in order for the gravitational coupling to be positive and the graviton to carry positive kinetic energy. Additionally, it must be $f^{\prime\prime}(R)>0$ to avoid local tachyonic instabilities [42, 43].
In the Friedmann-Lemaître-Robertson-Walker (FLRW) universe with line element in comoving polar coordinates $\left(t,r,\vartheta,\varphi\right)$
$ds^{2}=-dt^{2}+a^{2}(t)\left[\frac{dr^{2}}{1-kr^{2}}+r^{2}\left(d\vartheta^{2}+\sin^{2}\vartheta\,d\varphi^{2}\right)\right]$ (2.5)
(where $k=0,\pm 1$ is the normalized curvature index), the $f(R)$ field equations assume the form
$H^{2}=\frac{1}{3f^{\prime}}\left[8\pi\rho+\frac{Rf^{\prime}-f}{2}-3H\dot{R}f^{\prime\prime}\right]-\frac{k}{a^{2}}\,$,
$2\dot{H}+3H^{2}=-\frac{1}{f^{\prime}}\Big[8\pi P+(\dot{R})^{2}f^{\prime\prime\prime}+2H\dot{R}f^{\prime\prime}$
$+\ddot{R}f^{\prime\prime}+\frac{1}{2}\left(f-Rf^{\prime}\right)\Big]\,$, (2.7)
where $\rho$ and $P$ are the energy density and pressure of the cosmic matter fluid, respectively.
The evolution equation (LABEL:evolution_general2) for ${\cal K}{\cal T}$ in electrovacuum $f(R)$ gravity reads
$\frac{d\left({\cal K}{\cal T}\right)}{d\tau}={\cal K}{\cal T}\left(8\pi{\cal K}{\cal T}-\Theta\right)+\frac{2f(R)-Rf^{\prime}(R)}{24\pi f^{\prime}(R)}\,.$ (2.8)
In a FLRW universe, the gradient $\nabla^{a}\phi$ of the Brans-Dicke-like scalar field $\phi(t)=f^{\prime}(R(t))$ is always timelike. Denoting with an overdot the differentiation with respect to the comoving time $t$, we have
$\dot{\phi}=f^{\prime\prime}(R)\dot{R}$ (2.9)
and $\dot{\phi}$ has the sign of $\dot{R}$ since we require $f^{\prime\prime}(R)>0$. In order for
$u^{a}=\frac{\nabla^{a}\phi}{\sqrt{-\nabla^{c}\phi\nabla_{c}\phi}}=\frac{\nabla^{a}R}{\sqrt{-\nabla^{c}R\nabla_{c}R}}$ (2.10)
to be future-oriented, it must be $\dot{\phi}<0$ because
$\nabla^{a}\phi=g^{ab}\nabla_{b}\phi=g^{ab}\,{\delta_{b}}^{0}\,\partial_{t}\phi=-\dot{\phi}\,{\delta^{a}}_{0}$ (2.11)
and the time component $\nabla^{0}\phi=-\dot{\phi}$ must be positive for the effective $\phi$-fluid of our formalism to have $u^{0}>0$ and evolve toward the future. Therefore, according to Eq. (2.9), the thermal analogy is only applicable to FLRW solutions of $f(R)$ gravity for which
$R(t)=6\left(\frac{\ddot{a}}{a}+\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}\right)$ (2.12)
decreases with time.
In a FLRW universe, the proper time $\tau$ of the effective $\phi$-fluid coincides with the comoving time $t$. In fact, the four-velocity of the effective fluid has components
$u^{a}\equiv\frac{\nabla^{a}\phi}{\sqrt{-\nabla^{c}\phi\nabla_{c}\phi}}=\frac{-\dot{\phi}\,{\delta^{a}}_{0}}{|\dot{\phi}|}={\delta^{a}}_{0}$ (2.13)
in coordinates comoving with the effective $\phi$-fluid, therefore,
$u^{0}=\frac{dx^{0}}{d\tau}=\frac{dt}{d\tau}=1$ (2.14)
and $t=\tau$ up to an irrelevant additive constant corresponding to the choice of the origin on the time axis, so time evolution with respect to $t$ coincides with time evolution with respect to $\tau$.
In FLRW universes we have
${\cal K}{\cal T}=\frac{|\dot{\phi}|}{8\pi\phi}=\frac{f^{\prime\prime}(R)|\dot{R}|}{8\pi f^{\prime}(R)}\,.$ (2.15)
All solutions with $R=$ const. (when they exist) are automatically states of thermal equilibrium at ${\cal K}{\cal T}=0$ [44]. They include the radiation era with $R=0$ discussed in the next section.
Before proceeding to analyze the thermal view of $f(R)$ gravity, let us take a look at the phase space of spatially flat FLRW cosmology in these theories. The phase space of vacuum $k=0$ scalar-tensor cosmology has been described in [45]. When $k=0$ (the universe preferred by observations), the scale factor $a(t)$ enters the field equations (II), (2.7) only through the Hubble function $H\equiv\dot{a}/a$, which is a cosmological observable and is adopted as one of the phase space variables. The other variables are the scalar field $\phi$ and its time derivative $\dot{\phi}$. The phase space is, therefore, the 3D space $\left(H,\phi,\dot{\phi}\right)$, but the Hamiltonian constraint (a first order equation similar to an energy integral) forces the orbits of the solutions of the field equations to move on a 2D subset of this 3D space. The Hamiltonian constraint for vacuum, spatially flat, FLRW universes in Brans-Dicke gravity has the form [45]
$H^{2}=-H\,\frac{\dot{\phi}}{\phi}+\frac{\omega}{6}\,\left(\frac{\dot{\phi}}{\phi}\right)^{2}+\frac{V(\phi)}{6\phi}\,\,;$ (2.16)
solving this formal quadratic equation for $\dot{\phi}$ in terms of $H$ and $\phi$ yields [45]
$\dot{\phi}\left(H,\phi\right)=\frac{3\phi}{\omega}\left[H\pm\sqrt{H^{2}-\frac{2\omega}{3}\left(\frac{V}{6\phi}-H^{2}\right)}\,\right]\,.$ (2.17)
[…]
The structure of the 3D phase space of vacuum, $k=0$ $f(R)$ cosmology is simpler, and was discussed in [46]. Again, $a(t)$ enters the field equations only through $H$, and we choose the cosmological observable $\Theta=3H$ as one of the phase space variables. Since $f^{\prime}(R)>0$, there is a one-to-one correspondence between $R$ and $f(R)$ and we can choose the Ricci scalar $R$ as the second phase space variable. [^1] RR and $\Theta$ are independent because $R$ depends on both $\Theta$ and $\dot{\Theta}$ according to Eq. (2.12). Then, the third variable would be its derivative $\dot{R}$, but we can trade it for ${\cal K}{\cal T}=\frac{|\dot{\phi}|}{8\pi\phi}=\frac{f^{\prime\prime}(R)|\dot{R}|}{8\pi f^{\prime}(R)}$, limiting ourselves to the dynamical situations in which $\dot{R}<0$ to be able to apply the thermal formalism. Since $f(R)$ gravity corresponds to an $\omega=0$ Brans-Dicke theory (with a potential), the quadratic term in $\dot{\phi}/\phi$ is absent from the Hamiltonian constraint, which formally reduces to a linear equation, enabling us to eliminate $\dot{R}$ from the field equations using the single-valued function [46]
$\dot{R}\left(\Theta,R\right)$ $=$ $\frac{Rf^{\prime}(R)-f(R)-6H^{2}f^{\prime}(R)}{6Hf^{\prime\prime}(R)}$ (2.18)
$=$ $\frac{Rf^{\prime}(R)-f(R)-2\Theta^{2}f^{\prime}(R)/3}{2\Theta f^{\prime\prime}(R)}\,.$
Therefore, the orbits of the solutions are forced to live on the 2D subset of the 3D phase space $\left(\Theta,R,{\cal K}{\cal T}\right)$ described by
${\cal K}{\cal T}\left(\Theta,R\right)=\frac{1}{16\pi f^{\prime}(R)}\,\left|\frac{Rf^{\prime}(R)-f(R)-2\Theta^{2}f^{\prime}(R)/3}{\Theta}\right|\,.$ (2.19)
The projections of the orbits on the $\left(\Theta,R\right)$ plane cannot intersect, contrary to more general scalar-tensor theories in which two sheets project onto this plane [46].
[…]
$Rf^{\prime}(R)-f(R)+16\pi\Theta f^{\prime}(R){\cal K}{\cal T}<0\,.$ (2.20)
[…]
$\left(H,R,{\cal K}{\cal T}\right)=\left(\pm\sqrt{\frac{f_{0}}{6f_{0}^{\prime}}},\frac{2f_{0}}{f_{0}^{\prime}},0\right)$ (2.21)
(where $f_{0}\equiv f(R_{0})$ and $f^{\prime}_{0}\equiv f^{\prime}(R_{0})$). Since $k=0$, these fixed points can only be de Sitter spaces, possibly degenerating into the Minkowski point $\left(\Theta,R,{\cal K}{\cal T}\right)=\left(0,0,0\right)$. A radiation era with $R\equiv 0$ implies that also ${\cal K}{\cal T}\propto|\dot{R}|=0$ and its trajectory necessarily unfolds along the $\Theta$-axis, and is not a fixed point.
The 2D sheet on which the orbits live separates an “upper” and a “lower” region corresponding to $k=+1$ and k=1k=-1 (for $k\neq 0$, the orbits unfold in three phase space dimensions). These orbits cannot cross the $k=0$ “sheet” because the topology of 3-space cannot change dynamically.
III Thermal view of FLRW cosmology in general $f(R)$ gravity
Several results about the convergence of $f(R)$ gravity to GR (i.e., ${\cal K}{\cal T}\to 0$) or its departure from it (i.e., increasing ${\cal K}{\cal T}$) can be obtained for general functions $f(R)$. To begin with, the authors of [48] investigated the structural stability of the FLRW solutions of polynomial $f(R)$ gravity with respect to modifications of the field equations (from the Einstein equations to the fourth order equations (2.2) of polynomial $f(R)$ gravity). They studied early times near the initial singularity $t\to 0^{+}$ and late times $t\to+\infty$, assuming the matter source to be a perfect fluid with constant equation of state $P=w\rho$, with equation of state parameter in the range $1\leq w\leq 1$ (in particular for radiation $w=1/3$), and for all values of the curvature index $k$. This procedure makes sense only when the unperturbed solution is simultaneously a solution of GR and of $f(R)$ gravity. For general spacetimes, this is a very strong requirement and admits only a very restricted class of solutions [49, 50]. However, any viable theory of gravity should admit FLRW solutions, which are indeed very common. The structural stability of FLRW solutions of GR when the Einstein-Friedmann equations are modified to the fourth-order equations of $f(R)$ gravity was previously investigated by Barrow and Ottewill in the influential Ref. [47]. A notable result of this reference is that the FLRW solution for the radiation era solves exactly the fourth order equations of motion for any $f(R)$ gravity with $f(0)=0$ and $f^{\prime}(0)\neq 0$ (but only $f^{\prime}(0)>0$ is physically admissible).
To understand this result, which has remained a bit cryptic over the years, note that the class of $f(R)$ theories with $f(0)=0$ and $f^{\prime}(0)>0$ includes GR and its radiation solutions, which are (e.g., [51])
[…]
where the constant $\rho_{\mathrm{r}}^{(0)}$ is such that the energy density of radiation is $\rho_{\mathrm{r}}(a)=\rho_{\mathrm{r}}^{(0)}/a^{4}$. It is easy to see, using Eq. (2.12), that these solutions have Ricci scalar $R=0$ regardless of the form of the function $f(R)$, a feature that persists in all these $f(R)$ theories. In this radiation era, $R$ remains zero all the time and $\phi=f^{\prime}(R)=f^{\prime}(0)>0$ remains constant, that is, the theory is forced to be GR. From the thermal point of view, we have that ${\cal K}{\cal T}=0$ at all times for all these $f(R)$ theories, hence the structural stability of the solutions corresponds to the statement that the thermal state of equilibrium ${\cal K}{\cal T}=0$ is universal to these theories because $R$ is constant and the scalar degree of freedom $\phi=f^{\prime}(R)$ is simply eliminated from them. Time evolution in the $\left(\Theta,{\cal K}{\cal T}\right)$ plane is simple:
  • For $k=0$, $\Theta=\frac{3}{2t}$ is infinite at the Big Bang $t=0$ and vanishes as $t\to+\infty$. Time evolution is represented by motion strictly along the $\Theta$-axis, from $\Theta=+\infty$ towards $\Theta=0$.
  • For k=1k=-1, the universe begins with a Big Bang at $\Theta=+\infty$ and evolves towards the late-time $k=0$ solution with $\Theta=0$: the evolution in the $\left(\Theta,{\cal K}{\cal T}\right)$ plane is the same as in the previous case.
  • For $k=+1$, the universe begins with a Big Bang at $\Theta=+\infty$, expands to a maximum size where $\Theta=0$, and then contracts ($\Theta<0$) and ends in a Big Crunch at $\Theta\to$-$\infty$. Time evolution consists of motion along the ${\cal K}{\cal T}=0$ axis from $\Theta=+\infty$ to $\Theta=$-$\infty$.
Gravity never moves away from GR even though $\Theta$ becomes negative because it is always forced to stay on the $\Theta$-axis and to be GR, and the effective $\phi$-fluid disappears. In fact, the combination $\left(\nabla_{a}\nabla_{b}-g_{ab}\Box\right)\phi$ in Eq. (2.2) vanishes identically and this set of equations reduces to the Einstein equations.
From both the dynamical and the thermal points of view, as long as we restrict to radiation universes, these $f(R)$ theories are not interesting alternatives to GR because no propagating degree of freedom is added to the two (massless) tensor modes of GR, and there is never the possibility of gravity moving away from GR. Equation (2.8) is satisfied identically by ${\cal K}{\cal T}=0$.
Even with the restriction $f(0)=0$, the radiation solution (3.1) is not, in general, the only solution of the fourth order equations of $f(R)$ gravity and other solutions are possible (see the case of Starobinsky gravity below for an example).
The authors of [48] performed a perturbation analysis of GR solutions and found that in most (but not all) situations, the solutions of polynomial $f(R)$ gravity depart from those of GR (or, in their language are “non-perturbative”). They are singular solutions different from the GR universe with the same fluid content. This means that, in the $f(R)$ theories considered, gravity is significantly different from GR as $t\to 0^{+}$ unless special restrictions are imposed on the coefficients of the independent modes in linear perturbation theory [48].
The authors of [48] even conjectured that all “physically resonable” FLRW solutions of GR are unstable against structural perturbations of the field equations that change them from the Einstein equations to the field equations of higher-order gravity [48] (the radiation-dominated solutions of Barrow and Ottewill [47] already described constitute an obvious exception).
We can frame this conjecture within the thermal view of $f(R)$ gravity, using the recent results of [25]. The discussion is not limited to polynomial or other forms of the function $f(R)$, but is general.
First, it is crucial that the dynamics near a Big Bang singularity is dominated by gravity instead of matter. This behavior is quite plausible [52, 26], but is not a given and should be investigated case by case. When it turns out to be true, the dynamics reduces to that of vacuum $f(R)$ gravity.
We now know that, for vacuum or conformal matter, whether the conjecture of [48] is satisfied or not depends on the initial conditions. Expanding FLRW universes which start out with $8\pi{\cal K}{\cal T}>\Theta>0$, plus contracting ones with $\Theta<0$, depart from GR, but solutions that start out with $8\pi{\cal K}{\cal T}<\Theta$ do approach GR [25]. Solutions with $8\pi{\cal K}{\cal T}=\Theta$ and with constant ${\cal K}{\cal T}$ and constant $\Theta$ are fixed points, but they are thermally unstable [25]. The thermal picture settles the conjecture of [48] for vacuum or conformal matter with a completely different approach. For more general forms of matter, for which the term $\frac{T^{\mathrm{(m)}}}{(2\omega+3)\phi}$ in Eq. (LABEL:evolution_general2) is non-vanishing, the thermal picture is more complicated.
In their perturbative analysis, the authors of [53, 54, 48] find $k=0$ universes that are regular at $t=0$ (contrary to GR universes that have a Big Bang singularity) and converge to GR solutions as $t\to+\infty$, apparently contradicting the general thermal view of scalar-tensor gravity. The latter states that universes starting out with $8\pi{\cal K}{\cal T}>\Theta>0$ (which they do if they initially depart drastically from GR) run away from GR as long as $8\pi{\cal K}{\cal T}$ remains larger than $\Theta$ [25]. However, these universes of [53, 54, 48] are obtained for negative values of the parameter $\lambda^{2}\equiv\frac{f_{0}^{\prime}}{3f_{0}^{\prime\prime}}$ (the zero subscript denotes quantities evaluated on the unperturbed solution). Unbeknownst to the authors of Refs. [53, 54, 48] working in the 1970s and 1990s, we now know that in $f(R)$ gravity it must be $f^{\prime}(R)>0$ and $f^{\prime\prime}(R)>0$ [14, 15, 16], therefore, the region $\lambda^{2}<0$ is physically irrelevant. [^2] The same conclusion applies to the solutions (13), (15), (19), (21), and (27) of Ref. [48].
To continue the discussion, let us apply more results of Ref. [25] to the $f(R)$ subclass of Brans-Dicke theory. Gravity “heats up” if the effective potential (2.4) grows faster than $\sim\phi^{2}$ with $\phi=f^{\prime}(R)$, which is expressed by the asymptotic condition
$V(\phi)=Rf^{\prime}(R)-f(R)>\beta\left[f^{\prime}(R)\right]^{2}\,$, (3.2)
where $\beta$ is a positive constant. Using the modified notation $x\equiv R$ and $y(x)\equiv f(R)$, the non-linear ordinary differential equation setting the boundary of this situation,
$\beta y^{\prime 2}-xy^{\prime}+y=0\,$, (3.3)
admits the two linear solutions
$y_{1}(x)=\frac{c_{1}}{2\beta}\,x-\frac{c_{1}^{2}}{4\beta}\,,\quad\quad y_{2}(x)=-\frac{c_{1}}{4\beta}\,x-\frac{c_{1}^{2}}{16\beta}\,$, (3.4)
both of which give the boundary choice $f(R)=R-2\Lambda$ (i.e., GR) where, interestingly, only $\Lambda\geq 0$ is possible. This condition forbids, for example, Anti-de Sitter and asymptotically Anti-de Sitter universes. Hence, gravity “heats up” if $V(\phi)$ grows faster than $\phi^{2}$ or
$y^{\prime}\left(\beta y^{\prime}-x\right)<-y<0\,$, (3.5)
where $y>0$ and $y^{\prime}>0$. This condition is satisfied for $0<y^{\prime}<x/\beta$ (assuming $x\equiv R>0$), or $f^{\prime}(R)<R/\beta$, which means that $f(R)$ grows slower than $R^{2}/\left(2\beta\right)$. This condition on the derivative $f^{\prime}(R)$ can be translated into a condition on the function $f(R)$ if we require that $f(0)=0$. This condition (required in the Barrow-Ottewill theorem already discussed [47]) is satisfied by most physically relevant models in the literature, including $R^{n}$ and polynomial $f(R)$ gravity, as well as in popular $f(R)$ models aiming at explaining the cosmic acceleration without dark energy, such as the Hu-Sawicki scenario [55]
$f_{\mathrm{HS}}(R)=R-\frac{c_{1}m^{2}\left(R/m^{2}\right)^{n}}{c_{2}\left(R/m^{2}\right)^{n}+1}$ (3.6)
(where $c_{1,2}$, $m^{2}$, and $n$ are constants), the Starobinsky model [^3] Not to be confused with Starobinsky gravity $f(R)=R+\alpha R^{2}$. [56]
$f_{\mathrm{S}}(R)=R+\lambda R_{s}\left[\left(1+\frac{R^{2}}{R_{s}^{2}}\right)^{-q}-1\right]$ (3.7)
(with parameters $\lambda,R_{s},q$), and the Miranda-Joras-Waga-Quartin model [57]
$f_{\mathrm{MJWQ}}(R)=R-\alpha_{M}R_{*}\ln\left(1+\frac{R}{R_{*}}\right)$ (3.8)
(with two parameters $\alpha_{M},R_{*}$).
The condition $f(0)=0$ anchors all the relevant functions to the origin of the $\left(R,f(R)\right)$ plane and allows us to integrate the inequality $f^{\prime}(R)<R/\beta$ to $f(R)<R^{2}/(2\beta)$.
We conclude that:
  • gravity “heats up” if $f(R)$ grows slower than $R^{2}$ when $R$ increases (since we are limited to considering scenarios in which $\dot{R}<0$ in order for the formalism to apply, increasing time $t$ means decreasing $R(t)$ and vice versa);
  • gravity “cools” if $f(R)$ grows faster than $R^{2}$ when $R$ increases (i.e., $t$ decreases);
  • pure $R^{2}$ gravity (which is in some sense pathological [58, 59, 60, 61]), is the threshold between these two behaviors, corresponding to the quadratic potential $V(\phi)=\alpha\phi^{2}/2$, which disappears from the field equation (1.4) for $\phi$. Then the electrovacuum evolution equation for ${\cal K}{\cal T}$ reduces to
    $\frac{d\left({\cal K}{\cal T}\right)}{d\tau}={\cal K}{\cal T}\left(8\pi{\cal K}{\cal T}-\Theta\right)\,$, (3.9)
    which is the situation discussed in detail in Ref. [25].
In the context of braneworld models, similar conclusions have been reached for Starobinsky $f(R)=R+\alpha R^{2}$ gravity on the brane [62].
A comment is in order about the qualitative time evolution of $R(t)$, which coincides with that of $\phi(t)=f^{\prime}(R(t))$ since $f^{\prime}>0$. In order to apply the thermal formalism we need decreasing $R(t)$. Since $R(t)$ is continuous and monotonically decreasing, it either goes to a horizontal asymptote $R_{0}$ as $t\to+\infty$ (where $R_{0}\geq 0$ because the effective cosmological constant cannot be negative for the solutions (3.4), as already seen), or else $R(t)$ hits the $t$-axis from above with a non-zero derivative at a finite time $t_{0}$. Solutions with $R(t)\to+\infty$ as $t\to+\infty$ or to a finite time $t_{0}$ cannot be considered in the thermal formalism.
IV FLRW universes in polynomial and power-law $f(R)$ gravity
In this section, we consider polynomial functions
$f(R)=\sum_{n=1}^{m}a_{n}R^{n}$ (4.1)
(e.g., [47, 54, 48, 63, 64, 65]). $m=1$ corresponds to GR (possibly with a cosmological constant if we add a term $a_{0}\neq 0$ corresponding to $m=0$). In general, there is no point in adding a cosmological constant to these theories since one of the main motivations for $f(R)$ gravity is to explain away dark energy and the fine-tuned cosmological constant. The recent DESI results point against a cosmological constant anyway [66, 67, 68].
[…]
Conversely, if $R\to\infty$ approaching a singularity, the highest power of $R$ in the sum (4.1) dominates. An infinite $R$ could be present as an initial Big Bang singularity ($R(t)\to+\infty$ as $t\to 0^{+}$), [^4] However, one can still have a Big Bang singularity without $R$ diverging, as it happens in the radiation era where $R$ vanishes identically. as a final Big Crunch singularity ($R\to\infty$ as $t\to t_{0}^{-}$ with finite $t_{0}$), or as an asymptotic late-time singularity $R(t)\to$-$\infty$ as $t\to+\infty$. We do not discuss a Big Rip singularity with $R(t)\to+\infty$ as $t\to t_{0}$, where $t_{0}$ is finite, because $R(t)$ must be decreasing in order to apply the formalism.
There are many old results in the literature on scalar-tensor cosmology that are interesting for the thermal view and can be understood better and placed in a comprehensive framework using this new thermal formalism.
A body of work [53, 54, 47, 48, 63, 64] (partially summarized in [48]) studied FLRW cosmology in $f(R)$ gravity when the function $f(R)$ is polynomial. With the general polynomial form of $f(R)$, the potential $V(\phi)$ for the scalar degree of freedom $\phi$ is complicated and does not resemble scalar field potentials motivated by high energy physics. However, if $f(R)$ consists of a single power-law $f(R)=R^{n}$ (or if the highest power in the sum dominates for large $R$, or the lowest power dominates as $R\to 0$), then we know that the potential $V(\phi)=V_{0}\phi^{\beta}$ contributes to “cooling” gravity if $\beta<2$, equivalent to $n>2$ [25]. Vice-versa, gravity is “heated” if $\beta>2$, or $1<n<2$. In the case of purely quadratic gravity $f(R)=R^{2}$, which is a pathological theory in many respects [58, 59, 60, 61], the potential $V(\phi)$ is quadratic and is well known not to affect the equation of motion of $\phi$, which it enters only in the combination $\phi dV/d\phi-2V$. In this case, the evolution equation for ${\cal K}{\cal T}$ reduces to
$\frac{d\left({\cal K}{\cal T}\right)}{d\tau}={\cal K}{\cal T}\left(8\pi{\cal K}{\cal T}-\Theta\right)+\frac{T^{\mathrm{(m)}}}{3f^{\prime}(R)}$ (4.2)
and, for conformal matter with $T^{\mathrm{(m)}}=0$, it falls into the situation fully described in [25] (see Fig. 1), which applies to general scalar-tensor gravity and not only to FLRW cosmology. If spacetime begins with initial conditions such that $\Theta>0$ and $8\pi{\cal K}{\cal T}<\Theta$, it approaches GR. If the initial conditions are such that $\Theta<0$, or $8\pi{\cal K}{\cal T}>\Theta>0$, gravity always runs away from GR due to the scalar degree of freedom dominating over the GR degrees of freedom [25].
Figure 1: In the region of the $\left(\Theta,{\cal K}{\cal T}\right)$ plane between the half-line ${\cal K}{\cal T}=\Theta/(8\pi)$ (black, solid upper line) and the positive $\Theta$-axis, gravity “cools” and approaches GR. Outside of this region, gravity “heats up” and departs from GR. The red, dashed line corresponds to the solution (5.5)-(5.6) of vacuum $R^{n}$ gravity; the blue, dash-dotted line corresponds to the solution (5.15)-(5.17) of $R^{n}$ gravity with radiation. In both cases, $n=1.3$ and gravity moves towards the origin and GR.
For the function $f(R)$ given by Eq. (4.1), we have
$\phi$ $=$ $f^{\prime}(R)=\sum_{n=1}^{m}na_{n}R^{n-1}\,$, (4.3)
$V(\phi)$ $=$ $\sum_{n=2}^{m}(n-1)a_{n}R^{n}\,$, (4.4)
$V^{\prime}(\phi)$ $=$ $\frac{dV}{dR}\left(\frac{d\phi}{dR}\right)^{-1}=\frac{\sum_{n=2}^{m}n(n-1)a_{n}R^{n-1}}{\sum_{n=2}^{m}n(n-1)a_{n}R^{n-2}}=R\,.$
Since
$\phi V^{\prime}-2V=2f-\phi R=\sum_{n=1}^{m}(2-n)a_{n}R^{n}$ (4.6)
we have, for radiation or conformal matter with $T^{\mathrm{(m)}}=0$,
$\Box\phi$ $=$ $\frac{\phi V^{\prime}-2V}{3\phi}=\frac{1}{3}\left(\frac{2f}{\phi}-R\right)$ (4.7)
$=$ $\frac{\sum_{n=1}^{m}(2-n)a_{n}R^{n}}{3\sum_{n=1}^{m}\,na_{n}R^{n-1}}\,.$
The most important cases of polynomial $f(R)$ are power-law and Starobinsky gravity, which we examine next.
V Power-law $f(R)$
The case in which a single monomial $R^{n}$ appears in the polynomial $f(R)$ has been studied extensively in attempts to explain the present acceleration of the universe without dark energy [69, 70, 116, 71, 72, 73, 74, 75, 76, 16, 77, 78, 79, 80, 89, 90, 91, 92, 93, 94, 96, 82, 97, 95, 98, 99, 83, 100, 102, 103, 104, 81, 101, 105, 106, 107, 108, 109, 110, 84, 84, 85, 86, 112, 113, 114, 111, 87, 88, 115]. For $f(R)=R^{n}$, $f^{\prime}>0$ and $f^{\prime\prime}>0$ imply that $n>1$ (if $R>0$), while
$\phi=f^{\prime}(R)=nR^{n-1}\,,\quad\quad R=\left(\frac{\phi}{n}\right)^{\frac{1}{n-1}}\,$, (5.1)
and
$V(\phi)$ $=$ $Rf^{\prime}(R)-f(R)\Bigg|_{\phi=f^{\prime}(R)}=\left(n-1\right)R^{n}$ (5.2)
$=$ $\left(n-1\right)\left(\frac{\phi}{n}\right)^{\frac{n}{n-1}}$
or
$V(\phi)=V_{0}\,\phi^{\beta}\,,\quad\quad\beta=\frac{n}{n-1}$ (5.3)
with
$V_{0}=\frac{n-1}{n^{\frac{n}{n-1}}}\,.$ (5.4)
If $\beta<2$, corresponding to $n>2$, the potential terms “cool” gravity for $R>0$. If, instead, $\beta>2$ (equivalent to $1<n<2$), they “heat up” gravity for $R>0$.
Let us look now at two examples using exact FLRW solutions of the field equations of Brans-Dicke gravity which are also solutions of $R^{n}$ gravity. The first exact solution [116] corresponds to vacuum, $k=0$, a power-law potential $V(\phi)=V_{0}\phi^{\beta}$ and has power-law scale factor and scalar field
$a(t)$ $=$ $a_{0}\,t^{\frac{1+2\omega+\beta}{\left(2-\beta\right)\left(1-\beta\right)}}\equiv a_{0}\,t^{p}\,$, (5.5)
$\phi(t)$ $=$ $\phi_{0}\,t^{\frac{2}{1-\beta}}\,$, (5.6)
where $a_{0}$ and $\phi_{0}$ are positive constants. If we set $\omega=0$, this is also a solution of $R^{n}$ gravity with $\beta=\frac{n}{n-1}$ [116], provided that $n>1$ to keep $f^{\prime\prime}(R)>0$ (which also guarantees that $f^{\prime}(R)>0$). This universe expands (i.e., $p\equiv\frac{(2n-1)(n-1)}{2-n}>0$) if $n<1/2$ or if $1<n<2$. The relevant range is, therefore, $1<n<2$. The scalar field derivative
$\dot{\phi}=-\,\frac{2\phi_{0}\left(n-1\right)}{t^{2n-1}}$ (5.7)
is negative and so is $\dot{R}$, since
$H=\frac{p}{t}\,,\quad R(t)=\frac{6p\left(2p-1\right)}{t^{2}}\,.$ (5.8)
$R$ is positive when $p>1/2$, which corresponds to $5/4<n<2$, to which we restrict. We have
${\cal K}{\cal T}=\frac{n-1}{4\pi t}\to 0\quad\mbox{as}\;t\to+\infty$ (5.9)
and
$\Theta=3H=\frac{3(2n-1)(n-1)}{(2-n)t}\,$, (5.10)
so that
$8\pi{\cal K}{\cal T}=\frac{2(2-n)}{3(2n-1)}\,\Theta\,.$ (5.11)
We want to check whether the inequality $8\pi{\cal K}{\cal T}<\Theta$ is satisfied, which predicts the convergence to GR ${\cal K}{\cal T}\to 0$. Indeed, for the coefficient of $\Theta$, $\frac{2(2-n)}{3(2n-1)}$, to be larger than unity it must be $n<7/8$, which is never satisfied in the selected range $5/4<n<2$, therefore, the thermal formalism predicts that this solution converges to GR at late times, which it does (Fig. 1).
Let us look now at another exact solution of Brans-Dicke gravity in the presence of a perfect fluid with constant barotropic equation of state $P=\left(\gamma-1\right)\rho$, $V=V_{0}\phi^{\beta}$, $k=0$, and [116]
$a(t)$ $=$ $a_{0}\,t^{\frac{2\beta}{3\gamma\left(\beta-1\right)}}\,$, (5.12)
$\phi(t)$ $=$ $\phi_{0}\,t^{\frac{2}{1-\beta}}\,$, (5.13)
$\rho$ $=$ $\rho_{0}\,t^{\frac{2\beta}{1-\beta}}\,$, (5.14)
with constant $a_{0},\phi_{0},\rho_{0}$. Again, this is a solution of $R^{n}$ gravity if $\omega=0$ and $\beta=\frac{n}{n-1}$ with $n>1$, and we choose a radiation fluid ($\gamma=4/3$) for which $T^{\mathrm{(m)}}=0$, obtaining
$a(t)$ $=$ $a_{0}\,t^{n/2}\,$, (5.15)
$\phi(t)$ $=$ $\phi_{0}\,t^{2(1-n)}\,$, (5.16)
$\rho$ $=$ $\rho_{0}\,t^{-2n}\,.$ (5.17)
For this solution, $H=n/(2t)$, $R=\frac{3n(n-1)}{t^{2}}$ and $\dot{R}<0$, therefore the thermal formalism is applicable. We have
$8\pi{\cal K}{\cal T}=\frac{2(n-1)}{t}\to 0\,,\quad\Theta=\frac{3n}{2t}\,$, (5.18)
therefore,
$8\pi{\cal K}{\cal T}=\frac{4(n-1)}{3n}\Theta$ (5.19)
and the coefficient $\frac{4(n-1)}{3n}$ is less than unity for $1<n<4$. $\Theta$ starts infinite at $t=0$ and keeps decreasing, reaching zero as $t\to+\infty$. The point representing gravity in the $\left(\Theta,{\cal K}{\cal T}\right)$ plane moves along the straight line (5.19) towards the origin (Fig. 1).
To understand this situation in detail, we look at the evolution equation for ${\cal K}{\cal T}$
$\frac{d\left({\cal K}{\cal T}\right)}{dt}$ $=$ $8\pi\left({\cal K}{\cal T}\right)^{2}-\Theta{\cal K}{\cal T}+\frac{1}{24\pi}\left(V^{\prime}-\frac{2V}{\phi}\right)$
$=$ $8\pi\left({\cal K}{\cal T}\right)^{2}-\Theta{\cal K}{\cal T}+\frac{V_{0}(2-n)}{24\pi(n-1)}\,\phi^{\frac{1}{n-1}}\,.$
The left-hand side is $d\left({\cal K}{\cal T}\right)/dt=-\frac{(n-1)}{4\pi t^{2}}$, while the right- hand side results from the competition of cooling and heating terms:
${\cal K}{\cal T}\left(8\pi{\cal K}{\cal T}-\Theta\right)+\frac{V_{0}}{24\pi}\left(\frac{2-n}{n-1}\right)n^{\frac{1}{n-1}}\,R$
$=$ $\frac{(n-1)(n-4)}{8\pi t^{2}}+\frac{(2-n)(n-1)}{8\pi t^{2}}=-\frac{(n-1)}{4\pi t^{2}}$
which, of course, equals the left-hand side and shows how the heating term (e.g., for $1<n<2$)
$I\equiv 8\pi\left({\cal K}{\cal T}\right)^{2}+\frac{(2-n)(n-1)}{8\pi t^{2}}=\frac{(n-1)(3n-2)}{8\pi t^{2}}$ (5.22)
is smaller than the cooling term
$II\equiv\frac{-3n(n-1)}{8\pi t^{2}}\,.$ (5.23)
In fact, the ratio
$\left|\frac{I}{II}\right|=\left|\frac{3n-2}{3n}\right|$ (5.24)
is never larger than unity so the “cooling” term prevails for $1<n<2$.
V.1 Pure $R^{2}$-gravity
The case $n=2$ of pure quadratic gravity $f(R)=R^{2}$ corresponds to $\beta=2$ and to a quadratic potential for the effective Brans-Dicke scalar. It is also the strong gravity regime of Starobinsky gravity discussed in the next section, described by $f(R)=R+\alpha R^{2}\simeq\alpha R^{2}$ for large curvatures. In this special case the potential disappears from the equation of motion (1.4) of $\phi$ and from Eq. (LABEL:evolution_general2) and has no direct heating or cooling effect on gravity (cf. Eq. (LABEL:evolution_general2)).
We have $\Box\phi=-\left(\ddot{\phi}+3H\dot{\phi}\right)=0$, which admits the first integral
$\dot{\phi}=\frac{C}{a^{3}}$ (5.25)
well known in scalar-tensor cosmology. Since it must be $\dot{\phi}<0$ in order to apply the formalism, we restrict to values $C<0$ of the integration constant. On the other hand, $\phi=2R$ must be positive and the combined requirements $\phi>0$, $\dot{\phi}<0$ give that $R>0$ must always decrease. Therefore, as $t\to+\infty$, $R$ approaches a horizontal asymptote $R(t)\to R_{\infty}^{+}$, or else $R$ vanishes at a finite time $t_{0}$.
If $R_{\infty}>0$, we have an asymptotic de Sitter equilibrium state analogous to GR, but de Sitter spaces in $R^{2}$ gravity are unstable with respect to both homogeneous and inhomogeneous perturbations (Appendix A), so this situation is physically irrelevant.
If $R_{\infty}=0$, we have an asymptotic Minkowski space corresponding to $\phi=2R\simeq 0$ and infinite effective gravitational coupling $G_{\mathrm{eff}}\simeq 1/\phi$, which is clearly away from GR. Indeed, it is well known that $R^{2}$ gravity does not admit a Newtonian limit around Minkowski space [58] (although it does admit one around de Sitter space [117]). Similarly, a Hamiltonian analysis reveals the absence of propagating degrees of freedom around Minkowski space, or any background with $R=0$, in the linear approximation (although the full theory propagates three degrees of freedom) [118]. Counting the propagating degrees of freedom is far from trivial and depends on the background [119, 120, 121].
In principle, the remaining possibility is the “hard landing” in which $R$ vanishes at a finite time $t_{0}$ with $\dot{R}<0$: in this case
${\cal K}{\cal T}=\frac{|\dot{R}|}{8\pi R}\to+\infty$ (5.26)
and $R^{2}$ gravity runs infinitely far from GR.
VI Starobinsky gravity
Quadratic terms in the Lagrangian density are introduced by quantum corrections to GR, as in Starobinsky inflation [10],
$f(R)=R+\alpha\,R^{2}$ (6.1)
with $\alpha>0$ to guarantee $f^{\prime\prime}>0$ and where $R>-1/(2\alpha)$ to keep $f^{\prime}>0$. The trace equation (2.3) reduces to
$\Box R-\frac{R}{6\alpha}=\frac{4\pi}{3\alpha}\,T^{\mathrm{(m)}}\,.$ (6.2)
The other possible quadratic curvature corrections $R_{ab}R^{ab}$ and $R_{abcd}R^{abcd}$ are contained implicitly in the Starobinsky action because, in four spacetime dimensions (to which we restrict), the Gauss-Bonnet term $R^{2}-4R_{ab}R^{ab}+R_{abcd}R^{abcd}$ is a topological invariant,
$\int d^{4}x\,\sqrt{-g}\left(R^{2}-4R_{ab}R^{ab}+R_{abcd}R^{abcd}\right)=\mbox{const.}$ (6.3)
and, additionally, because in FLRW spaces
$\int d^{4}x\,\sqrt{-g}\left(R^{2}-3R_{ab}R^{ab}\right)=\mbox{const.}$ (6.4)
(e.g., [122]). Using Eqs. (6.3) and (6.4), one can trade $R_{abcd}R^{abcd}$ and $R_{ab}R^{ab}$ for terms in $R^{2}$ in the quadratic gravity action
$S=\int d^{4}x\,\sqrt{-g}\left(R+\alpha R^{2}+\beta R_{ab}R^{ab}+\gamma R_{abcd}R^{abcd}\right)\,.$ (6.5)
Quadratic gravity has been studied more extensively than other $f(R)$ theories. The Brans-Dicke theory equivalent to $f(R)=R+\alpha\,R^{2}$ has $\omega=0$,
$\phi=1+2\alpha R\,,\quad\quad R=\frac{1}{2\alpha}\left(\phi-1\right)\,$, (6.6)
and
$V(\phi)=\alpha R^{2}=\frac{1}{4\alpha}\left(\phi-1\right)^{2}\,.$ (6.7)
There are no de Sitter spaces in Starobinsky gravity (Appendix A), hence the states ${\cal K}{\cal T}=$ const. can only be Minkowski spaces ($k=0$), Anti-de Sitter spaces (k=1k=-1), or static Einstein universes ($k=+1$), if they exist, and they are physically relevant only if they are stable. [^5] Thermal states of equilibrium with ${\cal K}{\cal T}=$ const.>0>0 are, in principle, possible in scalar-tensor gravity but, thus far, they have been shown to be unstable or unphysical [123, 44, 124, 125]).
Assuming that the scalar degree of freedom dominates the cosmic dynamics, matter can be neglected and one can restrict to the vacuum field equations. The following discussion does not change if one considers a radiation fluid or other form of conformal matter with $T^{\mathrm{(m)}}=0$ (as done in [48]). Then, the equation for ${\cal K}{\cal T}$ becomes
$\frac{d\left({\cal K}{\cal T}\right)}{d\tau}={\cal K}{\cal T}\left(8\pi{\cal K}{\cal T}-\Theta\right)+\frac{\left(1-1/\phi\right)}{48\pi\alpha}\,$, (6.8)
or
$\frac{d\left({\cal K}{\cal T}\right)}{d\tau}={\cal K}{\cal T}\left(8\pi{\cal K}{\cal T}-\Theta\right)+\frac{R}{24\pi\left(1+2\alpha R\right)}\,$, (6.9)
while
${\cal K}{\cal T}\left(\Theta,R\right)=\frac{\left|\alpha R^{2}-2\left(1+2\alpha R\right)\Theta^{2}/3\right|}{16\pi\left(1+2\alpha R\right)|\Theta|}\,.$ (6.10)
This surface in the $\left(\Theta,R,{\cal K}{\cal T}\right)$ space is plotted in Fig. 2.
Figure 2: The orbits of the solutions of Starobinsky gravity are forced to lie on the surface ${\cal K}{\cal T}\left(\Theta,R\right)$ illustrated here in the 3D phase space $\left(\Theta,R,{\cal K}{\cal T}\right)$ (for illustration, we use $\alpha=10^{-3}$ and arbitrary units).
Since the scalar field potential (6.7) has an absolute minimum at $\phi=1$, at late times the dynamics is attracted to this stable fixed point, which corresponds to $R\to 0$. In other words, as time progresses, the universe expands, $R$ decreases (making it possible to apply the thermal formalism), vanishing asymptotically as $t\to+\infty$. In this limit, the quadratic term in $f(R)=R+\alpha R^{2}$ becomes irrelevant and the theory reduces to GR (without cosmological constant) with $f(R)\simeq R$. In the thermal point of view, ${\cal K}{\cal T}\to 0^{+}$.
Let us discuss the late-time limit $R\to 0$ and the strong gravity regime $R\to\infty$.
VI.1 Limit $R\to 0$
If $R\to 0$, the scalar field degree of freedom freezes, i.e., $\phi=1+2\alpha R\to 1$ and the potential disappears from the evolution equation of ${\cal K}{\cal T}$:
$\frac{d\left({\cal K}{\cal T}\right)}{d\tau}\simeq{\cal K}{\cal T}\left(8\pi{\cal K}{\cal T}-\Theta\right)\,.$ (6.11)
This situation is described in [25] and in Fig. 1) for general scalar-tensor gravity. Since $\dot{R}<0$, the limit $R\to 0$ implies that the axis $R=0$ is a horizontal asymptote for the function $R(t)$, with $\ddot{R}>0$.
At small curvatures $R\to 0$, the linear (GR) term dominates over the quadratic one in $f(R)=R+\alpha R^{2}$,
${\cal K}{\cal T}=\frac{|\dot{\phi}|}{8\pi\phi}=\frac{\alpha|\dot{R}|}{4\pi\left(1+2\alpha R\right)}\simeq\frac{\alpha|\dot{R}|}{4\pi}\to 0\,$, (6.12)
and GR is approached as $R(t)$ approaches its horizontal asymptote $R=0$. At this stage, if $R\to 0$ and assuming that it stays small, Ref. [25] and Fig. 1 tell us that initial conditions such that the dynamics begins below the half-line $8\pi{\cal K}{\cal T}=\Theta>0$ guarantee that gravity converges to GR while, if the universe begins above this line, or contracts with $\Theta<0$, it should run away from GR. So why does Starobinsky gravity always converge to GR as the universe expands? The key is that, as $R\to 0$, Eq. (6.10) yields
${\cal K}{\cal T}\left(\Theta,R\right)\simeq\frac{|\Theta|}{24\pi}\to 0\quad\mbox{as}\;\>\Theta\to 0$ (6.13)
hence $8\pi{\cal K}{\cal T}\simeq|\Theta|/3<\Theta$ and, in the late-time limit of large expansion $R\to 0$, the point representing Starobinsky gravity in the $\left(\Theta,{\cal K}{\cal T}\right)$ plane always lies below the critical half-line $8\pi{\cal K}{\cal T}=\Theta$, so gravity is bound to converge to GR.
VI.2 Strong gravity regime $R\to\infty$
In the strong gravity regime $R\to\infty$, $\phi\simeq 2\alpha R\to\infty$ and
$\frac{d\left({\cal K}{\cal T}\right)}{d\tau}\simeq{\cal K}{\cal T}\left(8\pi{\cal KT}-\Theta\right)+\frac{1}{48\pi\alpha}\,.$ (6.14)
The positive constant term in the right-hand side of Eq. (6.14) is a constant heat source which contributes to “heating” gravity and, if it was alone, it would cause ${\cal K}{\cal T}$ to grow linearly in time. It is obvious that for contracting universes ($\Theta<0$), the right-hand side is positive and $d({\cal K}{\cal T})/d\tau>0$, hence gravity departs from GR. However, this is not the only situation when this happens.
The constant curvature states with ${\cal K}{\cal T}=0$ are the only states with ${\cal K}{\cal T}=$ const. In fact, setting $d\left({\cal K}{\cal T}\right)/d\tau=0$ (where $d\left({\cal K}{\cal T}\right)/d\tau\equiv\psi\left(\Theta,{\cal K}{\cal T}\right)$) yields the formal algebraic equation
$\psi\left({\cal K}{\cal T},\Theta\right)\equiv 8\pi\left({\cal K}{\cal T}\right)^{2}-\Theta{\cal K}{\cal T}+\frac{1}{48\pi\alpha}=0$ (6.15)
with formal roots [^6] Equation (6.16) generalizes the equation $8\pi{\cal K}{\cal T}-\Theta=0$ used in the corresponding analysis of thermal equilibrium states ${\cal K}{\cal T}=$ const. when $\Box\phi=0$ [25].
$({\cal K}{\cal T})_{\pm}(\Theta)=\frac{1}{16\pi}\left(\Theta\pm\sqrt{\Theta^{2}-\frac{2}{3\alpha}}\,\right)\geq 0$ (6.16)
(which, if real, are non-negative if $\Theta>0$). They depend, of course, on the coefficient $\Theta$ appearing in Eq. (6.15). Obviously, ${\cal K}{\cal T}$ can only be constant if this coefficient $\Theta=3H$ is constant, which only allows for de Sitter or Anti-de Sitter spaces. Minkowski space and the Einstein static universe, which both have $\Theta=0$, are ruled out since they make the argument of the square root in the right-hand side of (6.16) negative, while there are no de Sitter spaces in this theory (Appendix A).
The function $\psi\left(\Theta,{\cal K}{\cal T}\right)$ is negative for
$\left({\cal K}{\cal T}\right)_{-}(\Theta)<{\cal K}{\cal T}<\left({\cal K}{\cal T}\right)_{+}(\Theta)\,$, (6.17)
therefore, ${\cal K}{\cal T}$ decreases and gravity “cools” in the region of the $\left(\Theta,{\cal K}{\cal T}\right)$ plane comprised between the curves $\left({\cal K}{\cal T}\right)_{\pm}(\Theta)$ given by Eq. (6.16), while it “heats up” in the other two regions ${\cal K}{\cal T}<({\cal K}{\cal T})_{-}$ and ${\cal K}{\cal T}>({\cal K}{\cal T})_{+}$ of this plane (see Fig. 3).
Instead of considering the curves $\left({\cal K}{\cal T}\right)_{\pm}(\Theta)$, one could consider the inverse form
$\Theta\left({\cal K}{\cal T}\right)=8\pi{\cal K}{\cal T}+\frac{1}{48\pi\alpha{\cal K}{\cal T}}\,.$ (6.18)
The curves $\left({\cal K}{\cal T}\right)_{\pm}(\Theta)$ themselves are not trajectories of the system. Next, one wonders whether single points on these curves can be fixed points during time evolution. To answer, for such a point to be a fixed point it must be $\Theta=$ const., while $d\left({\cal K}{\cal T}\right)/dt=0$ by virtue of this point being on the curves $\left({\cal K}{\cal T}\right)_{\pm}(\Theta)$. Then, the Ricci scalar
$R$ $=$ $6\left(\dot{H}+2H^{2}+\frac{k}{a^{2}}\right)=6\left(\frac{\dot{\Theta}}{3}+\frac{2\Theta^{2}}{9}+\frac{k}{a^{2}}\right)$ (6.19)
$=$ $\frac{4\Theta^{2}}{3}+\frac{6k}{a^{2}}$
is positive if $k=0,+1$ and a necessary (but not sufficient) condition for it to be negative is k=1k=-1, i.e., that this point describes Anti-de Sitter space. However, this statement would contradict the regime under discussion because $\phi=1+2\alpha R\simeq 2\alpha R$ when $R\to\infty$. Since $R<0$ for Anti-de Sitter space, the basic requirement that $\phi>0$ is violated. Therefore, Anti-de Sitter spaces on the curves ${\cal K}{\cal T}_{\pm}(\Theta)$ cannot be fixed points of the evolution and can, in principle, be crossed as time goes by.
Figure 3: In the region of the $\left(\Theta,{\cal K}{\cal T}\right)$ plane between the graphs of the functions ${\cal K}{\cal T}_{+}(\Theta)$ (solid) and ${\cal K}{\cal T}_{-}(\Theta)$ (dashed), gravity “cools” and approaches GR. Outside of this region, gravity “heats up” and departs from GR (here we set $\alpha=1$ for illustration).
In quadratic gravity there are spatially flat, radiation-dominated ($T^{\mathrm{(m)}}=0$) FLRW universes that avoid the initial Big Bang singularity and diverge in the future, $R(t)\to\infty$ as $t\to+\infty$ [53, 126, 127, 47]. Clearly, as $R\to\infty$ it is $f(R)=R+\alpha R^{2}\simeq\alpha R^{2}$ and Starobinsky gravity departs drastically from GR (but the perturbative approach leading to Starobinsky gravity will break down at some point in this limit). These solutions are found to be unstable and, therefore, unphysical [47]. In any case, they do not begin with $R=0$, but with a finite $R$. For these solutions, if $R(0)$ is larger than a critical value $R_{*}$, the solutions diverge as $t\to+\infty$, that is they diverge from GR. If they begin with $R(0)<R_{*}$, then $R(t)\to 0^{+}$ as $t\to+\infty$, approaching GR. The thermal view can be applied to these solutions with decreasing $R(t)$ and their behavior matches the results of the thermal view of scalar-tensor gravity [25] (applied to the specific subclass of Starobinsky gravity) and could have been predicted had this formalism been available in 1971. The thermal formalism cannot be applied to the solutions with $\dot{R}>0$.
Let us discuss now two other interesting classes of solutions of Starobinsky cosmology found numerically by Nariai and Tomita [126]. They studied the quadratic theory with action
$S=\int d^{4}x\sqrt{-g}\,\left[R+\eta\left(R^{2}+\alpha R_{ab}R^{ab}\right)\right]+S^{\textrm{(m)}}\,$, (6.20)
where matter is a radiation fluid, restricting themselves to $k=0$ FLRW solutions. Therefore, Eq. (6.4) applies and, on FLRW solutions, the theory can be recast as Starobinsky gravity $f(R)=R+\bar{\alpha}R^{2}$, where $\bar{\alpha}=\eta\left(1+\alpha/3\right)$. Nariai and Tomita looked for bouncing universes that avoid the Big Bang singularity [126]. Clearly, this is non-GR behavior because the radiation fluid always obeys the energy conditions and, in GR, a bounce with $\dot{H}>0$ requires the violation of the energy condition since $\dot{H}=-4\pi\left(\rho+P\right)$. This physics can be provided by the effective $\phi$-fluid which, being a fictitious fluid, can violate all of the energy conditions applicable to a real fluid. Nariai and Tomita integrated numerically the field equation for the scale factor $a(t)$ and found two types of solutions, which both begin by contracting ($\Theta=3H<0$) and reach an instant of time (which we will call $t=0$) at which $\dot{a}=0$ and $\Theta=3\dot{a}/a=0$ [126]. At times $t>0$, one class of solutions expands forever, asymptoting to the GR solution $a(t)\simeq\sqrt{t}$ as $t\to+\infty$, which means that ${\cal K}{\cal T}\to 0^{+}$ while $R\to 0$. The second class of solutions contracts again at $t>0$, ending in a Big Crunch with $\Theta\to$-$\infty$ and $R\to\infty$ at a finite time $t_{0}>0$ [126]. This means that $\ddot{a}$ changes from positive during the first contraction, to zero at $t=0$, to negative for $0<t<t_{0}$. This second type of solutions clearly departs from GR, but the thermal formalism is not applicable. In fact, for large $|R|$ it is $\phi\simeq 2\alpha R$, which implies that it must be $R>0$. At the finite time singularity $R$ diverges so it must be $R\to+\infty$. However, the Nariai-Tomita second class of solutions has $\dot{a}=0$ and $\ddot{a}=0$ at $t=0$, hence $R$ increases for times $0<t<t_{0}$, hence $\dot{R}<0$ is not satisfied, $\dot{\phi}>0$, and the thermal formalism is not applicable.
In the $\left(\Theta,{\cal K}{\cal T}\right)$ plane, these solutions of Starobinsky cosmology begin at negative $\Theta$ departing from GR and with upward-facing concavity, $\ddot{a}>0$. In this region, $R=6\left(\frac{\ddot{a}}{a}+\frac{\dot{a}^{2}}{a^{2}}\right)$ is positive and Eq. (6.9) becomes
$\frac{d\left({\cal K}{\cal T}\right)}{d\tau}={\cal K}{\cal T}\left(8\pi{\cal K}{\cal T}+|\Theta|\right)+\frac{|R|}{24\pi\left(1+2\alpha|R|\right)}\,;$ (6.21)
the right-hand side is positive and ${\cal K}{\cal T}$ can only increase departing from GR. When the solution reaches the bounce at $\Theta=0$ (a strongly non-GR feature), the solutions bifurcate: the first class enters a region where GR is approached and remains there, asymptotically approaching the GR solution $a(t)=a_{0}\sqrt{t}$. The second class of solutions goes back to the region $\Theta<0$, ending in a Big Crunch as $\Theta$ goes towards -$\infty$, but the formalism loses meaning.
Another result of quadratic gravity is that, if its spatially flat, radiation-dominated FLRW universes possess a Big Bang singularity, they approach the FLRW universes of GR as $t\to+\infty$ [47]. This result is generalized by the Barrow-Ottewill theorem [47] because Starobinsky gravity satisfies the hypothesis of the theorem $f(0)=0$ and $f^{\prime}(0)\neq 0$.
VII Conclusions
The thermal view of scalar-tensor gravity is ultimately only an analogy between this class of alternative theories of gravity and heat dissipation in a viscous fluid. However, like all analogies, there are two sides to it and one can learn much about gravity from physical intuition of the other side. Physical insight has been obtained in first-generation and Horndeski gravities, and here we have applied the thermal analogy specifically to $f(R)$ gravity, which is very popular to explain the current acceleration of the universe, provides the first scenario of inflation ever proposed (by Starobinsky), which is very successful, and is motivated by quadratic corrections to GR.
The main value of the thermal view of scalar-tensor gravity consists of its unifying power. The analysis of the previous sections has framed known results, some of which are old and remained rather cryptic, in the thermal view of $f(R)$ gravity and cosmology. These results remained disconnected for decades, and now fall into a single, coherent picture. The fourth order field equations of $f(R)$ gravity are rather complicated, even for spatially homogeneous and isotropic FLRW universes, and there is a large variety of situations corresponding to different matter contents of the universe, forms of the function $f(R)$, and initial conditions. It is logical, therefore, that one cannot catch all these possible theories and scenarios in a few single sentences. We find significant, however, that the new thermal view of $f(R)$ gravity unifies many apparently fragmented, and often forgotten, results, as discussed in the previous sections.
The results concerning scalar-tensor and $f(R)$ universes that we discussed here can be obtained by studying the dynamics ruled by the master equations of these theories (i.e., Eqs. 2.2), (2.3)) and, mathematically, do not require the thermal formalism (i.e., Eq. (2.8)) to be deduced. However, in many situations one would not think of investigating certain directions unless guided by the thermal view, and one often lacks a physical interpretation of these results without the unifying thermal framework.
Here we focussed on FLRW cosmologies and on expanding universes and we did not delve into the features of static or contracting universes, which are less interesting from the physical point of view, and will be presented elsewhere. There are also several results in the literature on spatially anisotropic Bianchi universes in $f(R)$ and in more general Horndeski cosmology, which seem relevant for the thermal view of scalar-tensor cosmology and will be analyzed separately.
Acknowledgements.
This work is supported, in part, by the Natural Sciences & Engineering Research Council of Canada (grant 2023-03234 to V. F.).
Appendix A Existence and stability of de Sitter spaces in vacuum $f(R)$ gravity
The condition for the existence of de Sitter spaces with Hubble function $H_{0}=$ const. and Ricci scalar $R_{0}=12H_{0}^{2}$ in vacuum $f(R)$ gravity is derived by substituting the de Sitter metric into Eq. (2.2), obtaining [47, 128]
$R_{0}f^{\prime}_{0}=2f_{0}\,$, (A.1)
where $f_{0}\equiv f(R_{0})$ and $f^{\prime}_{0}\equiv f^{\prime}(R_{0})$. If such a de Sitter space exists, the condition for its stability with respect to both homogeneous and inhomogeneus perturbations (to first order) is [128]
$\left(f^{\prime}_{0}\right)^{2}-2f_{0}f_{0}^{\prime\prime}\geq 0\,.$ (A.2)
For Starobinsky gravity with $f(R)=R+\alpha R^{2}$, the existence condition (A.1) cannot be satisfied and there are no de Sitter solutions in this theory.
In $R^{n}$ gravity, the existence condition for de Sitter spaces can be satisfied only if $n=2$. In this case, however, the stability condition becomes $R^{2}\geq 0$, which cannot be satisfied except for the degenerate Minkowski solution with $H_{0}=0$ and $R_{0}=0$ (a trivial de Sitter space).